# Regularized Langevin Dynamics for Combinatorial Optimization

### v1

@@ 4.1 Experimental Setup
Following Zhang et al. (2023), we use the Revised Model B (RB) graphs (Xu & Li, 2000) for the evaluation of MIS and MCl problems, and use the Barabasi-Albert (BA) graphs (Barabási & Albert, 1999) for the evaluation of MCut problem. In addition, we also include Erdős-Rényi (ER) graphs (Erdos & Rényi, 1984) used by Qiu et al. (2022) on the MIS problem. We follow the above works to generate RB, BA and ER graphs at two different scales. On RB and BA graphs, the small scale contains 200 to 300 nodes and the large scale contains 800 to 1200 nodes. While the ER graphs have a small scale of 700 to 800 nodes and a large scale of 9,000 to 11,000 nodes. The large-scale ER graphs are used for transfer testing on the models trained on the small-scale ER graph. A suffix of ‘-[$n$$N$]’ is used to differentiate the graphs with different scales, implying that the graphs contain $n$ to $N$ nodes. The test set size is 500 for RB and BA graphs, 128 for ER-[700700700700800800800800] and 16 for ER-[900090009000900011000110001100011000]. A training set of size 1000 and validation set of 500 graphs is used for all datasets except for ER-[900090009000900011000110001100011000]. The node weight in MIS and MCl is set as $1$ for all nodes.
@@
Following Qiu et al. (2022) and Zhang et al. (2023), we categorize our baselines as the classical operation research methods (OR), heuristic methods (H), reinforcement learning-based solvers (RL), supervised learning-based solvers (SL) and unsupervised learning-based solvers (UL). For MIS, we have the integer linear programming solver Gurobi (Gurobi Optimization, LLC, 2023) and MIS-specific solver KAMIS (Großmann et al., 2023) as the OR baselines, and the recent SA method iSCO (Sun et al., 2023) as a heuristic baseline. In the RL category, we include PPO (Ahn et al., 2020) and DIMES (Qiu et al., 2022). In the SL category, we have INTEL (Li et al., 2018b), DGL (Böther et al., 2022), and DIFUSCO (Sun & Yang, 2023). In the UL category, we use LTFT (Zhang et al., 2023) and DiffUCO (Sanokowski et al., 2024). For the non-MIS problems, the baselines include two OR methods, which are Gurobi and a semi-definite programming method (SDP); three heuristic methods, which are greedy, mean-filed annealing (MFA) (Bilbro et al., 1988) and iSCO (Sun et al., 2023); and three UL methods, which are ERDOES (Erdos & Rényi, 1984), LTFT (Zhang et al., 2023) and DiffUCO (Sanokowski et al., 2024), respectively. For most of those methods, we report their published results in Qiu et al. (2022); Zhang et al. (2023); Sun & Yang (2023); Sanokowski et al. (2024); Li et al. (2024). If a NN-based method has multiple variants, we only compare with the version with the longest running time (typically corresponding to the best result). For method iSCO, we run its code on our datasets as the time measurement in this work is inconsistent with others. For a fair comparison, we run iSCO with the same number of steps and trials as we did for RLSA.
@@
We used two servers for the training of RLNN, one with 8 NVIDIA RTX A6000 GPUs, and the other one with 10 NVIDIA RTX 2080 Ti GPUs. All the time measurement is conducted on a single A6000 GPU. We find the efficiency of RLNN highly susceptible to the inductive bias of the NN architecture, e.g., a two-parameter linear model is enough to fit the gradient of MIS in Equation 15. Therefore, we mainly focus on verifying the effectiveness of RLNN algorithm, without optimizing the neural network architecture. In our experiment, we parameterized RLNN with a five-layer GCN (Kipf & Welling, 2017) with 128 hidden dimensions. Due to the increasing computational complexity at each step, we also accordingly reduce the number of sampling steps and trials of RLNN compared to RLSA, with other hyperparameters kept the same. We include more details, such as the hyperparameters, in Appendix A
Figure 1: Comparison between RLD and standard discrete LD for SA. RLD/RLSA is in the red color. The man value of the primal gap in the test set is plotted, while the shaded area indicates the standard deviation.
@@ 4.2 Main Results
In performance evaluation, we compare the mean value of the achieved problem-specific objective (larger is better) of each method on each problem, including the set size for MIS, clique size for MCl and cut size for MCut. In addition, we compare the total running time (lower is better) of each method on the entire test set by sequentially evaluating each instance. Since the OR solvers are guaranteed to find the optimal solution with enough running time, we do not include them for comparison.
Table 1 presents the results on the MIS problem. It can be seen that with a similar or even less running time, RLSA shows a significant improvement against the SOTA NN-based methods on RB and ER graphs. Besides, RLSA also consistently outperforms another gradient-guided SA method iSCO, with the same number of steps and trials. Due to the simplicity of RLSA, it only take around 5-20% running time of iSCO, but with a better objective value. The performance of RLNN is also competitive, taking the second-best result on 2 out of 4 datasets. It also achieves the similar performance on ER-[90090090090011000110001100011000] to iSCO, using less than 20% running time. Due to the significant increase in computational overhead at each step, RLNN only uses 10% number of trails and a shorter sampling chain compared to RLSA. We observe that RLNN could sometimes outperform RLSA with the same number of running steps and trials, which indicates the potential of NN in learning the problem distribution. But how to balance the computational resources between the per-step computation and searching efforts would an interesting topic.
The comparative results on the MCl and MCut problems are summarized in Table 2. Generally, RLSA still maintains a clear efficiency advantage against iSCO. We observe that iSCO has a similar performance with RLSA on MCl at both scales, but RLSA and RLNN are still clearly better than other baselines. RLSA takes a consistent lead on MCut, while RLNN, DiffUCO and iSCO have a similar performance, with a significant improvement against other baselines.
Besides, we also include the comparative results between iSCO and RLSA with 10×10$\times$10 × running steps in Appendix B, and the observation is consistent. In general, we find both RLSA and RLNN competitive on our evaluation benchmarks, and RLSA shows an impressive performance on all datasets with only a limited computational cost.
Table 3: Ablation Study on regularization in RLNN. The numbers correspond to the set size (larger is better).
@@ 4.3 Ablation Study
To verify the effectiveness of regularization in RLD, we conduct the ablation study on RLSA and RLNN, respectively. We first compare RLD with the standard discrete LD (Zhang et al., 2022) for SA, searching the step size $\alpha$ over the set $\{0.1,0.01,0.001\}$. All other hyperparameters are kept the same as in RLSA. Figure 1 compares the dynamics of the primal gap (Berthold, 2014) across the sampling process. Here, the primal gap on each instance is defined as
@@
It can be seen that the standard discrete LD always ends up at a sub-optimal solution except on MCl. The searching could easily get stuck in a local optima, indicated by the flat stage. In contrast, RLSA usually converges in just a few steps (less than 100) without getting stuck in any local optima. Note that our searching set has already covered the most common choices of the gradient descent step size for continuous data, and a smaller step size, e.g., 0.001, shows an even worse performance. Contrast to the optimization in the continuous domain, CO clearly presents a different challenge and we address it by RLD.
We then examine the regularization term in Equation 20 for training RLNN. We perform the ablation study on small-scale graphs by training RLNN without regularization, and present the comparative results in Table 3. The regularization brings a significant improvement on RLNN in most cases, except on MCut. We hypothesize this is because MCut has no constraint and suffers less from the local optima. But on other benchmarks, we find it almost impossible to effectively train RLNN without regularization. Here we visualize training dynamics on MIS and MCl by plotting the set/clique size on the validation set (larger is better) in Figure 2.
@@
It can be seen that the performance of the orange curve (no regularization) remains unchanged (on MIS) or even drops after more training epochs (on MCl). The usage of a local optimization loss function makes the diversity of training samples a huge concern. While the regularization would enforce RLNN to collect different training samples and encourage the search during inference time. The similar spirit is also used in some well-known reinforcement learning algorithms to encourage the exploration, such as the curiosity (Pathak et al., 2017) and soft policies (Haarnoja et al., 2017, 2018).
5 Related Work
5.1 Neural Solvers for Combinatorial Optimization
The neural network (NN) models have recently garnered vast attention in solving CO problem (Bengio et al., 2020). The NN-based solvers could be roughly categorized into three classes according to the training methods, including the supervised learning-based (Li et al., 2018a; Gasse et al., 2019; Sun & Yang, 2023; Li et al., 2023, 2024), unsupervised learning-based (Karalias & Loukas, 2020a; Wang et al., 2022; Wang & Li, 2023; Zhang et al., 2023; Sanokowski et al., 2024), and reinforcement learning-based (Khalil et al., 2017; Qiu et al., 2022) methods. Our proposed RLNN method is partially based on reinforcement learning, but could be efficiently trained with a local objective. Such a feature has greatly improved its training efficiency by eliminating the need to estimate the future return.
5.2 Sampling for Combinatorial Optimization
Sampling-based methods (Metropolis et al., 1953; Hastings, 1970; Neal, 1996; IBA, 2001) have been commonly applied in CO problems (WANG et al., 2009; Bhattacharya et al., 2014; Tavakkoli-Moghaddam et al., 2007; Seçkiner & Kurt, 2007; Chen & Ke, 2004). However, earlier methods often encountered slow convergence compared to learning-based approaches due to an inefficient proposal. Recent advancements in discrete Monte Carlo Markov Chain (Grathwohl et al., 2021; Zhang et al., 2022; Sun et al., 2022) have revitalized sampling-based methods and Sun et al. (2023) demonstrated that simulated annealing (SA) can surpass neural CO solvers. In this work, we have advanced the current study on discrete Langevin dynamics, and proposed a novel SA algorithm. Our conclusion supports the previous study the further advances the development of the field.
@@ 6 Conclusion & Limitation
In this work, we point out the difference in local optimal between continuous optimization and continuous optimization (CO), and propose a novel sampling framework called Regularized Langevin Dynamics (RLD) to tackle the issue in CO. On top of that, we develop two CO solvers, one based on simulated annealing (SA), and the other one based on neural networks. Our empirical evaluation on three classical CO problems demonstrate that our proposed methods can achieve the state-of-the-art (SOTA) or near-SOTA performance with high efficiency. Especially, our proposed SA method consistently outperforms the previous SA baseline using only 20% running time. In summary, RLD is a simple yet effective framework, showing a strong potential in addressing CO problems.
In this work, we only consider binary data for ease of analysis. Although the whole framework could be generalized, its effectiveness remains unclear on other CO problems with integer or mixed integer variables. Future work may extend it to more real-world CO problems by taking other conditions into consideration. Besides, we have given an intuitive explanation of RLD in this work, but the theoretical understanding of RLD is generally missing. We also expect to address this part in our future work.
@@ Appendix A Additional Experiment Details
A.1 Hyperparameters
We include the hyperparameters of RLSA and RLNN in Table 4 and 5, respectively. The initial temperature $\tau_{0}$ is randomly searched in the range of 0.001 to 10, and the step size is randomly searched in the range of 2 to 100. In general, larger $K$ and $T$ always lead to a better performance, so we simply control them such that the running time of RLSA is similar to the fastest baseline. For RLNN, we allow for more time budget, just controlling its $K$ and $T$ such that its running time is comparable to most baselines.
@@ A.1 Hyperparameters
A.2 Training
RLNN is parameterized by a five-layer GCN (Welling & Teh, 2011) with 128 hidden dimensions. A linear layer is first used to project the input solution $\mathbf{x}$ into a 128-dim embedding $\mathbf{H}^{0}$. Each layer of GCN performs the following update:
$\mathbf{H}^{l+1}=\sigma(\mathbf{U}^{l}\mathbf{H}^{l}+\mathbf{V}^{l}\mathbf{D}^{-1/2}\hat{\mathbf{A}}\mathbf{D}^{-1/2}\mathbf{H}^{l})+\mathbf{H}^{l},$ (24)
where $\mathbf{U}^{l}$ and $\mathbf{V}^{l}$ are model parameters, $\hat{\mathbf{A}}=\mathbf{A}+\mathbf{I}_{N\times N}$ is the adjacency matrix with the self loop, $\mathbf{D}$ is a diagonal degree matrix with $\mathbf{D}_{ii}=\sum{j=1}^{N}\hat{\mathbf{A}}_{ij}$, and $\sigma$ is the activation function. We use ReLU as the activation function for all layers. The output hidden vector is projected by a final linear layer into the single dimension. A sigmoid function is then applied to generate the flipping probability $q_{\theta}(\mathbf{x}^{\prime}_{i}=1-\mathbf{x}_{i}|\mathbf{x})$.
We train RLNN with $50$ epochs in all data sets except RB- [700700700700800800800800] for MCl, where we use 80 epochs since we notice that the training does not converge at 50 epochs. On each graph, we sample $K^{\prime}$ trajectories with $T^{\prime}$ lengths, which amounts to $K^{\prime}T^{\prime}$ samples to train on. The batch size is set as 32 per GPU, and an Adam optimizer is used for optimization, with a learning rate of 0.0001.
In terms of the training time, all experiments can finish under 1 hour on our server with 8 RTX A6000 GPUs. The large-scale graphs and ER-[$700$-$800$] graphs may need 3 hours to finish on our server with 10 RTX 2080 Ti GPUs. Our implementation is based on PyTorch Geometric (Fey & Lenssen, 2019) and accelerate (Gugger et al., 2022).
During our inference time, we use the float16 data type to store the tensor, which accelerates the tensor product. But in general, we find our method way more efficient than our baselines even without this technique.

### v2

@@ 5.1 Experimental Setup
Following Zhang et al. (2023), we evaluate MIS and MCl using Revised Model B (RB) graphs (Xu & Li, 2000), and we evaluate MCut using Barabási-Albert (BA) graphs (Barabási & Albert, 1999). Following Qiu et al. (2022), we also include Erdős-Rényi (ER) graphs for MIS. As in prior work, each graph type is generated at two scales:
  • RB and BA graphs: small (200200200200300300300300 nodes) and large (8008008008001200120012001200 nodes).
  • ER graphs: small (700700700700800800800800 nodes) and large (900090009000900011000110001100011000 nodes).
The large-scale ER graphs serve as a transfer‐testing set for models trained on small-scale ER. We append a suffix “-[ $n$$N$ ]” to each graph name to indicate its size range.
For RB and BA (both scales) and ER-[700700700700800800800800], we use 1000 graphs for training, 100 for validation, and 500 (RB/BA) or 128 (ER-[700700700700800800800800]) for testing. ER-[900090009000900011000110001100011000] is reserved solely for testing (16 graphs).
@@
Following Qiu et al. (2022) and Zhang et al. (2023), we categorize our baselines as the classical operations research solvers (OR), (human-designed) heuristic solvers (H), supervised learning-based solvers (SL), reinforcement learning-based solvers (RL), and unsupervised learning-based solvers (UL). For MIS, we have the integer linear programming solver Gurobi (Gurobi Optimization, LLC, 2023) and MIS-specific solver KaMIS (Großmann et al., 2023) as the OR baselines, and the recent SA method iSCO (Sun et al., 2023) as a heuristic baseline. In the SL category, we include INTEL (Li et al., 2018b), DGL (Böther et al., 2022), and DIFUSCO (Sun & Yang, 2023). In the RL category, we have PPO (Ahn et al., 2020) and DIMES (Qiu et al., 2022). In the UL category, we use LTFT (Zhang et al., 2023) and DiffUCO (Sanokowski et al., 2024). For the non-MIS problems, the baselines include two OR methods, which are Gurobi and a semi-definite programming method (SDP) for MCut; three heuristic methods, which are greedy, mean-field annealing (MFA) (Bilbro et al., 1988) and iSCO (Sun et al., 2023); and three UL methods, which are EGN (Karalias & Loukas, 2020), LTFT (Zhang et al., 2023) and DiffUCO (Sanokowski et al., 2024), respectively. For most of these methods, we report their published results (Qiu et al., 2022; Zhang et al., 2023; Sun & Yang, 2023; Sanokowski et al., 2024; Li et al., 2024). If an NN solver has multiple variants, we only compare with the variant with the longest runtime (typically corresponding to the best result). We rerun the official code of iSCO [^1] https://github.com/google-research/discs on our datasets due to the inconsistency of time measurement in this work, where the average runtime of iSCO is compared to the total runtime of its baselines. For a fair comparison, we run iSCO with the same number of steps and trials as we do for RLSA.
@@
Our implementation is based on PyTorch Geometric (Fey & Lenssen, 2019) and Accelerate (Gugger et al., 2022). We use two servers for the RLNN training, one with 8 NVIDIA RTX A6000 GPUs and the other with 10 NVIDIA RTX 2080 Ti GPUs. We evaluate all methods on the first server, using a single A6000 GPU. We find the efficiency of RLNN highly susceptible to the inductive bias of the NN architecture, e.g., a two-parameter linear model is enough to fit the gradient of MIS in Equation 19. Since the main focus of our evaluation is to verify the effectiveness of RLD under the approximate gradient, we keep the model architecture of RLNN basically the same as the ones used in prior works (Qiu et al., 2022; Sanokowski et al., 2024). Future work may further optimize the neural architecture by injecting more prior knowledge about the problem structure, as in Yau et al. (2024). In our experiment, we parameterize RLNN with a five-layer GCN (Kipf & Welling, 2017) with 128 hidden dimensions. Due to the increasing computational complexity at each step, we also accordingly reduce the number of sampling steps and trials of RLNN compared to RLSA, with other hyperparameters kept the same. We include more details in Appendix A.
Figure 1: Primal‐gap trajectories for SA solvers using RLD versus the standard discrete LD method (Zhang et al., 2022). The RLD (corresponding to RLSA) curve is shown in red, and the remaining curves (in distinct colors) correspond to discrete LD with various step‐size settings. Solid lines denote the mean primal gap over the test set, and shaded regions represent the standard deviation.
@@ 5.2 Main Results
In performance evaluation, we compare the mean value of the achieved problem-specific objective (larger is better) of each method on each problem, including the set size for MIS, clique size for MCl and cut size for MCut. In addition, we compare the total runtime (lower is better) of each method throughout the test set by sequentially evaluating each instance. Since OR solvers are guaranteed to find the optimal solution with enough runtime, we do not include them for comparison.
Table 1 reports our results on the MIS problem. RLSA achieves significant improvements over SOTA NN‐based methods on both RB and ER graphs, with similar or even shorter runtime. Moreover, RLSA consistently outperforms iSCO, another gradient‐guided SA method, with the same number of sampling steps and independent trials. Due to its parallel sampling design, RLSA requires only 5-20% of iSCO’s runtime while delivering better objective values.
Based on less prior knowledge (using an approximate gradient), RLNN shows a lower performance than RLSA, but still claims the second-best result in two of the four datasets. On ER-[900090009000900011000110001100011000] graphs, RLNN matches iSCO’s performance but uses under 20% of the runtime. Compared to the SOTA NN baselines, RLNN shows a performance comparable to DiffUCO on the RB graphs and clearly outperforms DIMES on ER. Moreover, it should be pointed out that DiffUCO can require up to two days of training even on small graphs, e.g., RB-[200200200200300300300300], whereas RLNN completes the training in under three hours regardless the graph scales, while delivering comparable performance. This highlights RLNN’s superior training efficiency against previous diffusion models, thanks to its local training objective.
Table 2 summarizes our comparative results on MCl and MCut. RLSA retains a clear efficiency advantage over iSCO. Although iSCO matches RLSA’s performance on MCl at both scales, RLSA and RLNN still outperform all other baselines. On MCut, RLSA consistently leads; while RLNN, DiffUCO, and iSCO achieve comparable results—all significantly better than the remaining methods.
We also include extended comparisons between iSCO and RLSA with ten times more sampling steps in Appendix B, which confirms the above findings. Overall, both RLSA and RLNN remain highly competitive across our benchmarks, with RLSA delivering impressive results on every dataset at minimal computational cost.
Table 3: Ablation study on the effectiveness of regularization in RLNN. The numbers correspond to the set size (larger is better).
@@ 5.3 Ablation Study
To verify the effectiveness of the proposed regularization in RLD, we conduct the ablation study on RLSA and RLNN, respectively. We first compare RLD with the standard discrete LD (Zhang et al., 2022) for SA, searching the step size $\alpha$ over the set $\{0.1,0.01,0.001\}$. All other hyperparameters are kept the same as in RLSA. Figure 1 compares the dynamics of the primal gap (Berthold, 2014) across the sampling process. Here, the primal gap on each instance is defined as
@@
It is evident that the standard discrete LD always ends up at a sub-optimal solution except on MCl. The searching could easily get stuck in a local optimum, indicated by the flat stage. In contrast, RLSA typically converges in fewer than 100 steps without becoming trapped. Note that our search set already includes the most common gradient-descent step sizes used for continuous problems, and even a smaller step size (e.g., 0.001) yields worse results. Unlike continuous optimization, combinatorial optimization presents unique challenges; RLD is specifically designed to address these.
We next perform an ablation study on the regularization term in Equation 24 for RLNN training. Specifically, we train RLNN on small‐scale graphs both with and without regularization, and report results in Table 3. Adding regularization significantly improves RLNN’s performance in most cases, with the exception of MCut—likely because MCut is unconstrained and is less prone to local optima. On all other benchmarks, training RLNN without regularization proves almost ineffective. Figure 2 visualizes this effect by plotting the set/clique size on the validation set (larger is better) for MIS and MCl.
@@
From Figure 2, the curve without regularization (orange) remains flat on MIS or even deteriorates after more training epochs on MCl. This highlights that only using the unsupervised loss of EGN makes sample diversity a critical concern. By contrast, the regularization term encourages RLNN to gather more diverse training samples and explore more effectively during inference, resulting in the stable performance gains shown by the blue curve throughout training.
In fact, the idea of RLNN also parallels well‐known reinforcement learning techniques to encourage exploration, such as curiosity-driven exploration (Pathak et al., 2017) and soft policies (Haarnoja et al., 2017, 2018).
@@ 6 Conclusion & Limitation
In this work, we point out the specific challenge from the local optima in combinatorial optimization (CO), and propose a novel sampling framework called Regularized Langevin Dynamics (RLD) to tackle the issue in CO. On top of that, we develop two CO solvers, one based on simulated annealing (SA), and the other one based on neural networks. Our empirical evaluation on three classic CO problems demonstrate that our proposed methods can achieve state-of-the-art (SOTA) or near-SOTA performance with high efficiency. In particular, our proposed SA method consistently outperforms the previous SA baseline using only 20% running time, while RLNN significantly improves the training efficiency of previous diffusion models. In summary, RLD is a simple yet effective framework that shows great potential in addressing CO problems.
In this work, we only consider binary data for ease of analysis. Although the whole framework could be generalized, its effectiveness remains unclear in other CO problems with categorical, integer, or mixed integer variables. Future work may also extend it to other CO problems with global constraints, such as the Traveling Salesman Problem. Besides, we have only given an intuitive explanation of RLD in this work, but the theoretical understanding of RLD is generally missing. We also expect to address this part in the future.
@@ Appendix A Additional Experiment Details
A.1 Overview of Hyperparameters
Tables 4 and 5 summarize the hyperparameters for RLSA and RLNN, respectively. We use random search to determine the initial temperature $\tau_{0}$ within $[0.001,10]$ and the regularized step size $d$ within $[2,100]$. Since larger values of $K$ and $T$ generally yield better performance, we chose $K$ and $T$ for RLSA so that its runtime matches that of the fastest baseline on each dataset. For RLNN, we allow a larger time budget by selecting $K$ and $T$ such that its runtime is comparable to most baselines. The inference is performed using the float16 data type to accelerate tensor operations.
@@ A.1 Overview of Hyperparameters
A.2 Implementation of RLNN
RLNN is parameterized by a five-layer GCN (Kipf & Welling, 2017) with 128 hidden dimensions. A linear layer is first used to project the input $\mathbf{x}$ to a 128-dim embedding $\mathbf{H}^{0}$. Each layer of GCN performs the following update:
$\mathbf{H}^{l+1}=ReLU(\mathbf{U}^{l}\mathbf{H}^{l}+\mathbf{V}^{l}\mathbf{D}^{-1/2}\hat{\mathbf{A}}\mathbf{D}^{-1/2}\mathbf{H}^{l})+\mathbf{H}^{l},$ (28)
where $\mathbf{U}^{l}$ and $\mathbf{V}^{l}$ are the model parameters at $l$-th layer, $\hat{\mathbf{A}}=\mathbf{A}+\mathbf{I}_{N\times N}$ is the adjacency matrix with the self loop, $\mathbf{D}$ is a diagonal degree matrix with $\mathbf{D}_{ii}=\sum_{j=1}^{N}\hat{\mathbf{A}}_{ij}$. The output hidden representation is projected to a scalar via a linear layer, and then a sigmoid activation yields the flipping probability $q_{\theta}(\mathbf{x}^{\prime}_{i}=1-\mathbf{x}_{i}|\mathbf{x})$ for each node.
We train RLNN with 50 epochs on all datasets, except RB-[700700700700800800800800] for MCl, where we use $80$ epochs because the model does not converge in 50 epochs. On each graph, we sample $K^{\prime}$ trajectories of length $T^{\prime}$, yielding $K^{\prime}T^{\prime}$ training samples. The batch size is set to 32 per GPU, and we optimize with Adam at a learning rate of 0.0001.
In terms of training time, our model completes training on small-scale graphs (except ER-[700700700700800800800800]) in under half an hour using eight RTX A6000 GPUs, and the larger instances finish within one hour; on a server with ten RTX 2080 Ti GPUs, runtimes are slightly longer, but the longest experiment still finishes within three hours. By comparison, baseline methods such as DiffUCO (Sanokowski et al., 2024) require two days to train a single model on RB–[200200200200300300300300]. We attribute this efficiency advantage to the local training objective of RLNN, which avoids estimating long-term high-variance reward signals with standard reinforcement learning methods, as employed by DiffUCO.
Note that the float32 data type is used during RLNN training, which is switched to float16 at the inference time.